# HiLo: A Learning Framework for Generalized Category Discovery Robust to Domain Shifts

### v1

@@ D.1 Proof of bounds for the target error
From Theorem 3.4 of [81], we can know that:
@@
Consider an ideal joint hypothesis $g^{*}$, which is the hypothesis which minimizes the combined error (ideally zero). By using the triangle inequality [80], we have:
@@ D.2 A tighter bound for the target error
Using Pinsker’s Inequality [82], we have:
@@
As Equation 4 is Donsker-Varadhan representations of KL divergence [83] that approximates the mutual information using a neural network (MLP) $\Phi$, the bound on mutual information in terms of the JS divergence is:
$I(\boldsymbol{z}_{d};\boldsymbol{z}_{s})$ $=\mathcal{D}_{KL}(\mathcal{Z}^{d}\|\mathcal{Z}^{s})$
@@
Appendix E Incorporating UDA techniques for GCD with domain shifts
Given the large body of research in Unsupervised Domain Adaptation (UDA), here, we study of effectiveness of incorporating the state-of-the-art UDA techniques into the baseline methods and our HiLo method to augment them. Differently to our task, domain adaptation does not consider discovering categories in the unlabelled images from the unseen domain. Specifically, we consider incorporating MCC [84] and NWD [16] techniques for all the baseline methods to evaluate their performance on GCD with domain shifts. MCC can be adapted by minimizing between-class confusion and maximizing within-class confusion, while NWD uses a single task-specific classifier with implicit K-Lipschitz constraint to obtain better robustness for all the domain adaptation scenarios. Results are shown in Table 6. By comparing the results with those of ‘Real + Painting’ in Table 2 of the main paper, we can see that the results for each method are marginally improved by introducing these techniques. This reveals that simply adopting the UDA techniques to the GCD methods is not sufficient to handle the challenging problem of GCD with domain shifts. Moreover, we can see that HiLo again notably outperforms all other methods after introducing these UDA techniques, despite the gain by these techniques being relatively marginal, further demonstrating the effectiveness and significance of our HiLo design.
Real Painting
Methods All Old New All Old New
RankStats+ +MCC+NWD 37.3 62.1 23.4 31.0 51.2 9.2
UNO+ 46.9 72.4 32.8 32.1 47.6 17.7
ORCA 33.4 50.1 26.7 30.0 41.1 9.1
GCD 50.6 54.0 48.4 34.0 43.1 22.7
SimGCD 63.1 77.1 56.9 35.7 39.0 32.4
HiLo (Ours) 65.0 77.8 58.0 42.5 43.1 42.0
Furthermore, in Table 7 we extend our analysis by examining various UDA techniques (e.g., Mixstyle, EFDM, UniOT) and data augmentation methods (e.g., Mixup, Cutmix). The results reveal that while some UDA techniques and data augmentations offer improvements over the original SimGCD, they fall short of addressing the challenges of GCD with domain shifts. Specifically, Mixstyle slightly enhances SimGCD’s performance on both $\Omega^{a}$ and $\Omega^{b}$. EFDM improves SimGCD’s performance only on $\Omega^{a}$, likely due to its reliance on the assumptions that the unlabelled data are from the target domain only and the target domain is single, which are not the case for the challenging task of GCD with domain shifts. Mixup and Cutmix yield performance boost for SimGCD on both $\Omega^{a}$ and $\Omega^{b}$, but if we replace our proposed PatchMix constrastive learning in HiLo with them, the performance has a clear drop. Meanwhile, we also apply UniOT, a universal domain adaption method, on our task due to its architectural similarity with SimGCD, both utilizing a single classifier for known categories and a cluster head for novel ones. To enable UniOT for GCD with domain shifts, we modify the output dimensions of UniOT’s classifier and cluster head to match the total number of semantic categories in the unlabelled data. When using the default hyperparameters of UniOT, we found that it struggles to produce acceptable results. We then run a grid search to find the optimal hyperparameters for UniOT, during which we find the performance of UniOT on our task is highly sensitive to the choice of OT solver and its associated hyperparameters, restricting its reliability on our task. UniOT achieves slightly inferior results compared to HiLo on the seen domain, while showing significantly lower performance on the unseen domain. In summary, HiLo consistently outperforms all baselines enhanced with these UDA and data augmentation techniques. These results underscore the necessity of our tailored approach for GCD with domain shifts, demonstrating that simply adopting existing UDA or data augmentation methods is clearly insufficient for this challenging task.
Table 7: Influence of different UDA techniques (e.g., Mixstyle, EFDM) and data augmentations (e.g., Mixup, Cutmix). We also compare HiLo with one baseline method UniOT from universal domain adaptation.
Real Painting
Methods All Old New All Old New
Reference SimGCD 61.3 77.8 52.9 34.5 35.6 33.5
SimGCD + Mixstyle 62.3 76.8 54.0 35.0 36.1 34.0
SimGCD + EFDM 62.6 76.0 54.7 34.1 34.8 33.8
SimGCD + Mixup 62.7 76.5 54.3 34.9 37.2 32.5
SimGCD + Cutmix 62.5 76.3 54.1 33.2 36.0 31.6
Reference HiLo 64.4 77.6 57.5 42.1 42.9 41.3
HiLo + Mixup - PatchMix 63.7 77.1 56.8 39.8 39.9 38.7
HiLo + Cutmix - PatchMix 62.9 76.8 56.0 37.4 38.0 36.7
Reference HiLo 64.4 77.6 57.5 42.1 42.9 41.3
UniOT 63.3 77.4 57.4 35.3 38.7 32.0
Appendix F Effect of mutual information minimization on different datasets
@@
In Figure 7, we further analyze the effectiveness of the Mutual Information (MI) regularization by comparing the results with and without MI, on DomainNet and SSB-C. For experiments on DomainNet, the seen domain is the ‘Real’ domain, while for experiments on SSB-C the seen domain is the original SSB. The results demonstrate that MI regularization’s efficacy is particularly pronounced in two distinct scenarios: datasets with markedly distinct low-level styles w.r.t. the data from the seen domain (e.g., ‘Real’ vs. ‘Quickdraw’, ‘Real’ vs. ‘Infograph’) and those with closely related semantic categories within the dataset (e.g., each individual dataset in SSB-C contains instances closely related by a single shared semantic superclass). In the case of Quickdraw, the dramatic style variations are readily captured by low-level features, allowing MI regularization to effectively disentangle semantic features from low-level information. Note that models trained on Quickdraw exhibit unsatisfying performance, as samples are too abstract to align the semantics in supervised contrastive learning, making it difficult for the model to extract semantically discriminative features robust to domain shifts. For the SSB-C benchmark, where object structures remain largely similar across different noise types, the MI regularization plays a more significant role in distinguishing subtle semantic changes without being overly affected by the domain shifts.
@@
Figure 7: Comparison between HiLo with and without MI across different datasets. DomainNet is a generic dataset with disparate styles (e.g., Painting, Quickdraw, Sketch, Clipart and Infograph). The labelled data are all from ‘Real’ domain. SSB-C (e.g., CUB-C, Scars-C and FGVC-C) is created by adding several common noises in the real world to the fine-grained datasets. The labelled data are all from the original SSB. The improvement by MI regularization is pronounced when low-level styles are distinct w.r.t. the data from the seen domain (e.g., ‘Real’ vs. ‘Quickdraw’, ‘Real’ vs. ‘Infograph’) or when the semantics are closely related within the dataset (e.g., each individual dataset in SSB-C contains instances closely related by a single shared semantic superclass).
Appendix G Novel category discovery in the presence of domain shifts
@@
In this paper, we consider a challenging problem of generalized category discovery with domain shifts, which, to our knowledge, has not been studied in the literature. However, the study of novel category discovery under domain shifts has been considered in [51] from a domain adaptation perspective, which introduces a self-labeling framework, called NCDD, that can categorize unlabelled images from both source and target domains, by maximizing mutual information between labels and input images. The unlabelled images from the target domain may contain images from new categories that are not present in the source domain. Differently, in our study, we consider that unseen classes are also present in the unlabelled data from the source domain (i.e., the domain $\Omega^{a}$), and new domains may appear at test time. Meanwhile, our HiLo framework also differs significantly from NCDD. Particularly, HiLo learns to disentangle domain-semantic features by minimizing the mutual information between domain and semantic heads. It also incorporates a novel PatchMix contrastive learning method and a curriculum learning approach to facilitate the robustness of representation to domain shifts. To compare HiLo with NCDD, we reimplement the NCDD method [^2] Our NCDD reimplementation’s performance aligns with other efforts [85] and experiment on the experimental configuration following [51] on the CUB-C. We present the results in Table 8. As can be seen, HiLo significantly outperforms NCDD and all other baselines, highlighting its effectiveness on domain-semantic disentanglement.
@@
Appendix H Investigation of CLIP for GCD with domain shifts
CLIP [86] has demonstrated strong performance in various computer vision tasks. We thus investigate its potential for the challenging problem of GCD with domain shifts. We employ the pretrained vision transformer from CLIP as the backbone for HiLo. As illustrated in Table 9, employing CLIP significantly improves the performance of HiLo on DomainNet, compared with the DINO-based HiLo and SimGCD.
@@ Appendix H Investigation of CLIP for GCD with domain shifts
Table 9, verifies that a strong visual encoder can bring performance boost on both seen and novel domains. In Table 10, we further compare with another two CLIP baselines to better understand the potential of the visual language model, i.e., zero-shot CLIP with oracle class names (which are not expected to be unavailable in GCD) and with zero-shot CLIP a very large vocabulary (i.e., WordNet [87]), where we conduct zero-shot inference using the class names of both known and unknown classes, by comparing the visual feature of each image and the text features of class descriptions. The results in Table 10 demonstrate that CLIP models do not enhance robustness compared to our HiLo model, a visual-only model, on CUB-C, despite that an extra vocabulary is provided for the CLIP model (which arguably reduces the difficulty of the GCD task, in which we do not assume any extra textual or visual knowledge on the unlabelled data). This finding aligns with [88], which indicates that robustness under natural distribution shifts does not necessarily translate to robustness under synthetic distribution shifts, thereby suggesting the limited impact of CLIP models on covariate shifts. HiLo outperforms CLIP and CLIP, despite that they ‘cheat’ by using an extra vocabulary, which further underscores the effectiveness and robustness of HiLo and the challenge of GCD with domain shifts.
Table 10: Zero-shot performance of CLIP on CUB. CLIP is the CLIP with oracle class names, while CLIP is the CLIP with a large vocabulary (i.e., WordNet [87]).
@@
Appendix I Detailed evaluation of SSB-C datasets
@@
Appendix J Additional experimental results on DomainNet
@@
In Table 14, we report the results by considering the ‘Infograph’ domain as $\Omega^{b}$. ‘Others’ denotes the results on the unlabelled data from the remaining 4 domains aside from ‘Real’ and ‘Infograph’. Table 15 shows the evaluation on each domain in ‘Others’. Similarly, Table 16 and Table 17 show the results by considering ‘Quickdraw’ domain as $\Omega^{b}$; Table 18 and Table 19 show the results by considering ‘Sketch’ domain as $\Omega^{b}$; and Table 20 and Table 21 show the results by considering ‘Clipart’ domain as $\Omega^{b}$. Our HiLo framework consistently outperforms baseline methods in both ’All’ and ’New’ performance. Notably, the ’Quickdraw’ domain presents greater challenges than other domains due to its highly abstract and difficult-to-recognize images, resulting in unsatisfactory performance for all methods.
@@
Appendix K Clustering domain-specific and semantic-specific features
In Figure 8, we present the visualization by first applying PCA to the domain features and semantic features obtained through $\tilde{\mathcal{H}}$, and then plotting the corresponding images. As can be seen, the images are naturally clustered according to their domains and semantics, demonstrating that HiLo successfully learns domain-specific and semantic-specific features.
Figure 8: Visualization of domain and semantic features via projecting them through PCA. We randomly sample instances from the entire dataset and apply PCA to project the semantic and domain features into a 2-dimensional space. Our results demonstrate that the domain branch tends to cluster images based on covariate features, while the semantic branch clusters images based on categories. Best viewed in PDF with zoom.
Appendix L HiLo on the vanilla GCD setting
@@
Appendix M Effects of different output dimensions for the semantic and domain heads
@@
Appendix N Unknown category number
As the total number of semantic categories cannot be accessed in the real-world setting, we evaluate our HiLo with an estimated number of categories using an off-the-shelf method [8] on CUB (see Table 24). We find that our method consistently outperforms the strong baseline when the exact number of categories is unknown.
@@
Appendix O Hyperparameter choices for HiLo components
The hyperparameters of HiLo can be grouped via each component: (a) PatchMix (i.e., $\beta_{k}$); (b) representation learning and parametric classification losses (i.e., $\tau$, $\lambda$, $\epsilon$); (c) curriculum learning (i.e., $r_{0}$, $r^{\prime}$ and $t^{\prime}$). We follow [17, 5] to choose values for the shared hyperparameters in (a) and (b) respectively.
As summarized in Table 25, we choose the hyperparameters in (a) and (b) following [17] and [5] respectively. For the hyperparameters in (c), we choose the values through the validation split of the labelled data in the ‘Orignal’ domain.
@@
$\tau_{u}$ 0.07 Suggested values following [5]
$\tau_{c}$ 1.0 Suggested values following [5]
$\tau_{s}$ 0.1 Suggested values following [5]
$\tau_{t}$ 0.07 Suggested values following [5]
$\lambda$ 0.35 Suggested values following [5]
$\beta$ $\sim\operatorname{Beta}(\log(1+e),\log(1+e))$ Suggested value following [17]
@@
Appendix P Effects of learning rates
@@
Our study aims to extend AI systems’ capabilities from closed-world to open-world scenarios, particularly enhancing next-generation AI systems to categorize and organize open-world data autonomously. Despite promising results on public datasets, our method has limitations. First, interpretability needs improvement, as the underlying decision-making principles remain unclear. Second, cross-domain robustness is inadequate. Although our method has achieved the best overall and new class discovery results in the GCD setting with domain shifts, performance still has significant room for improvement. Third, the novel domains we investigated in the paper are still limited. Fourth, our current method is not specifically developed to handle imbalanced datasets, so the results may not be optimal without explicitly and jointly accounting for domain shifts and class imbalance.

### v2

@@ D.1 Proof of bounds for the target error
From Theorem 3.4 of Kifer et al. (2004), we can know that:
@@
Consider an ideal joint hypothesis $g^{*}$, which is the hypothesis which minimizes the combined error (ideally zero). By using the triangle inequality Ben-David et al. (2006), we have:
@@ D.2 A tighter bound for the target error
Using Pinsker’s Inequality Csiszár & Körner (2011), we have:
@@
As Equation 4 is Donsker-Varadhan representations of KL divergence Donsker & Varadhan (1983) that approximates the mutual information using a neural network (MLP) $\Phi$, the bound on mutual information in terms of the JS divergence is:
$I(\bm{z}_{d};\bm{z}_{s})$ $=\mathcal{D}_{KL}(\mathcal{Z}^{d}\|\mathcal{Z}^{s})$
@@
Appendix E Effect of mutual information minimization on different datasets
@@
Our comprehensive analysis of HiLo’s performance across diverse datasets, with and without Mutual Information (MI) regularization, is presented in Figure 6. The results demonstrate that MI regularization’s efficacy is particularly pronounced in two distinct scenarios: datasets with markedly distinct low-level styles (e.g., Quickdraw, Infograph) and those with closely related semantic categories (e.g., SSB-C benchmark). In the case of Quickdraw, the dramatic style variations are readily captured by low-level features, allowing MI regularization to effectively disentangle semantic features from low-level information. Note that Quickdraw samples are still too abstract to learn from constrastive learning, leading to poor performance of the model. For the SSB-C benchmark, where image structure remains largely unchanged across different noise types, MI regularization proves crucial in distinguishing subtle semantic differences from low-level information variations.
@@
Figure 6: Comparison between HiLo with and without MI across different datasets. DomainNet is a generic dataset with disparate styles (e.g., Painting, Quickdraw, Sketch, Clipart and Infograph). The labelled data are all from ‘Real’ domain. SSB-C (e.g., CUB-C, Scars-C and FGVC-C) is created by adding several common noises in the real world to the fine-grained datasets. The labelled data are all from the original SSB. When low-level style is quite different (e.g., Infograph, Quickdraw) or semantics are close (e.g., SSB-C benchmark), the improvement of MI regularization is pronounced.
Appendix F Novel category discovery in the presence of domain shifts
@@
In this paper, we consider a challenging problem of generalized category discovery with domain shifts, which, to our knowledge, has not been studied in the literature. However, the study of novel category discovery under domain shifts has been considered in Yu et al. (2022) from a domain adaptation perspective, which introduces a self-labeling framework, called NCDD, that can categorize unlabelled images from both source and target domains, by maximizing mutual information between labels and input images. The unlabelled images from the target domain may contain images from new categories that are not present in the source domain. Differently, in our study, we consider that unseen classes are also present in the unlabelled data from the source domain (i.e., the domain $\Omega^{a}$), and new domains may appear at test time. Meanwhile, our HiLo framework also differs significantly from NCDD. Particularly, HiLo learns to disentangle domain-semantic features by minimizing the mutual information between domain and semantic heads. It also incorporates a novel PatchMix contrastive learning method and a curriculum learning approach to facilitate the robustness of representation to domain shifts. To compare HiLo with NCDD, we reimplement the NCDD method [^1] Our NCDD reimplementation’s performance aligns with other efforts An et al. (2023) and experiment on the experimental configuration following Yu et al. (2022) on the CUB-C. We present the results in Table 8. As can be seen, HiLo significantly outperforms NCDD and all other baselines, highlighting its effectiveness on domain-semantic disentanglement.
@@
Appendix G Investigation of CLIP for GCD with domain shifts
CLIP Radford et al. (2021) has demonstrated strong performance in various computer vision tasks. We thus investigate its potential for the challenging problem of GCD with domain shifts. We employ the pretrained vision transformer from CLIP as the backbone for HiLo. As illustrated in Table 9, employing CLIP significantly improves the performance of HiLo on DomainNet, compared with the DINO-based HiLo and SimGCD.
@@ Appendix G Investigation of CLIP for GCD with domain shifts
Table 9, verifies that a strong visual encoder can bring performance boost on both seen and novel domains. In Table 10, we further compare with another two CLIP baselines to better understand the potential of the visual language model, i.e., zero-shot CLIP with oracle class names (which are not expected to be unavailable in GCD) and with zero-shot CLIP a very large vocabulary (i.e., WordNet Miller (1995)), where we conduct zero-shot inference using the class names of both known and unknown classes, by comparing the visual feature of each image and the text features of class descriptions. The results in Table 10 demonstrate that CLIP models do not enhance robustness compared to our HiLo model, a visual-only model, on CUB-C, despite that an extra vocabulary is provided for the CLIP model (which arguably reduces the difficulty of the GCD task, in which we do not assume any extra textual or visual knowledge on the unlabelled data). This finding aligns with Taori et al. (2020), which indicates that robustness under natural distribution shifts does not necessarily translate to robustness under synthetic distribution shifts, thereby suggesting the limited impact of CLIP models on covariate shifts. HiLo outperforms CLIP and CLIP, despite that they ‘cheat’ by using an extra vocabulary, which further underscores the effectiveness and robustness of HiLo and the challenge of GCD with domain shifts.
Table 10: Zero-shot performance of CLIP on CUB. CLIP is the CLIP with oracle class names, while CLIP is the CLIP with a large vocabulary (i.e., WordNet Miller (1995)).
@@
Appendix H Detailed evaluation of SSB-C datasets
@@
Appendix I Additional experimental results on DomainNet
@@
In Table 14, we report the results by considering the ‘Infograph’ domain as $\Omega^{b}$. ‘Others’ denotes the results on the unlabelled data from the remaining 4 domains aside from ‘Real’ and ‘Infograph’. Table 15 shows the evaluation on each domain in ‘Others’. Similarly, Table 16 and Table 17 show the results by considering ‘Quickdraw’ domain as $\Omega^{b}$; Table 18 and Table 19 show the results by considering ‘Sketch’ domain as $\Omega^{b}$; and Table 20 and Table 21 show the results by considering ‘Clipart’ domain as $\Omega^{b}$. Our HiLo framework consistently outperforms baseline methods in both ‘All’ and ‘New’ performance. Notably, the ’Quickdraw’ domain presents greater challenges than other domains due to its highly abstract and difficult-to-recognize images, resulting in unsatisfactory performance for all methods.
@@
Appendix J HiLo on the vanilla GCD setting
@@
Appendix K Effects of different output dimensions for the semantic and domain heads
@@
Appendix L Unknown category number
As the total number of semantic categories cannot be accessed in the real-world setting, we evaluate our HiLo with an estimated number of categories using an off-the-shelf method Vaze et al. (2022) on CUB (see Table 24). We find that our method consistently outperforms the strong baseline when the exact number of categories is unknown.
@@
Appendix M Hyperparameter choices for HiLo components
The hyperparameters of HiLo can be grouped via each component: (a) PatchMix (i.e., $\beta_{k}$); (b) representation learning and parametric classification losses (i.e., $\tau$, $\lambda$, $\epsilon$); (c) curriculum learning (i.e., $r_{0}$, $r^{\prime}$ and $t^{\prime}$). We follow Zhu et al. (2023); Wen et al. (2023) to choose values for the shared hyperparameters in (a) and (b) respectively.
As summarized in Table 25, we choose the hyperparameters in (a) and (b) following Zhu et al. (2023) and Wen et al. (2023) respectively. For the hyperparameters in (c), we choose the values through the validation split of the labelled data in the ‘Orignal’ domain.
@@
$\tau_{u}$ 0.07 Suggested values following Wen et al. (2023)
$\tau_{c}$ 1.0 Suggested values following Wen et al. (2023)
$\tau_{s}$ 0.1 Suggested values following Wen et al. (2023)
$\tau_{t}$ 0.07 Suggested values following Wen et al. (2023)
$\lambda$ 0.35 Suggested values following Wen et al. (2023)
$\beta$ $\sim\operatorname{Beta}(\log(1+e),\log(1+e))$ Suggested value following Zhu et al. (2023)
@@
Appendix N Effects of learning rates
@@
Appendix O Stability of different methods
As the differences in the results of the GCD benchmark tests can be very large, we obtain the averaged results in Table 2 and Table 3 by conducting three independent runs for each method on both DomainNet and SSB-C. Here we visualize the bar chart of ‘All’ classes ACC for each methods and list the corresponding error lines generated by the these independent runs. We notice that the error bars of ORCA and SimGCD exhibit significant oscillations.
Figure 8: The ‘All’ ACC results are averaged by conducting three independent runs for each method on both DomainNet and SSB-C.
Appendix P Qualitative Results
We provide qualitative results on DomainNet and CUB-C. In Figure 9, we present the visualization by first applying PCA to the domain features and semantic features obtained through $\tilde{\mathcal{H}}$, and then plotting the corresponding images. As can be seen, the images are naturally clustered according to their domains and semantics, demonstrating that HiLo successfully learns domain-specific and semantic-specific features.
Figure 9: Visualization of domain and semantic features via projecting them through PCA. We randomly sample instances from the entire dataset and apply PCA to project the semantic and domain features into a 2-dimensional space. The domain branch tends to cluster images based on covariate features, while the semantic branch clusters images based on categories. Best viewed in PDF with zoom.
The attention map offers valuable insights into the focus of Transformer-based models on the input. We obtain the attention maps for the CLS token from multiple attention heads in the final layer of the ViT backbone, highlighting the top 10% most attended patches in Figure 10. We observe that, compared with the baseline, HiLo is much more effective in focusing on the foreground object even in the presence of significant domain shifts (e.g., painting style, foggy weather). This demonstrates that HiLo is robust to domain shifts and remains unaffected by potential spurious correlations between semantic features and low-level statistics.
@@
Our study aims to extend AI systems’ capabilities from closed-world to open-world scenarios, particularly enhancing next-generation AI systems to categorize and organize open-world data autonomously. Despite promising results on public datasets, our method has limitations. First, interpretability needs improvement, as the underlying decision-making principles remain unclear. Second, cross-domain robustness is inadequate. Although our method has achieved the best overall and new class discovery results in the GCD setting with domain shifts, performance still has significant room for improvement. Third, the novel domains we investigated in the paper are still limited. Domain and class imbalance present additional challenges in GCD scenarios. Our current method was not specifically developed to handle these issues, which are important areas for future work.